OpenStudy (anonymous):

find the slope of the line that passes through each pair of points (-3,-2) (-3,2)

OpenStudy (anonymous):

m=y2-y1 over x2-x1 So m= 2-(-3) over -3(-3)

OpenStudy (anonymous):

so 5 over 0

OpenStudy (anonymous):

-3-(-3) is o